Find the instantaneous rate of change of s(t) = t^2 + 4 at t = 6.

Answer by stanbon(75887)

Find the instantaneous rate of change of s(t) = t^2 + 4 at t = 6.
----
Find the derivative:
ds/dt(t) = 2t
--
Evaluate at t = 6
ds/dt(6) = 2*6 = 12
=========================
Cheers,
Stan H.